# NASU – Novel Actuating Screw Unit: Origami-inspired Screw-based Propulsion on Mobile Ground Robots

### v2

@@ I Introduction
Figure 1: NASU is a screw-based propulsion unit for robot ground mobility that dynamically adjusts its angle of attack. The use of NASU is to allow control for optimizing for efficiency or velocity over a wide array of terrains. It allows for trading off higher traveling velocity and lower efficiency (left) for higher efficiency and lower velocity (right) for resource-constrained robotic deployments.
@@ I-B Related Works
Some of the first studies in screw propulsion and locomotion showed theoretically and experimentally that velocity tends to increase with increasing angle of attack, while thrust force decreases in sand and soil-like environments [18, 19]. Later computational and experimental analyses confirmed a higher thrust force is produced with smaller angles of attack in the case of granular media composed of silica beads [25, 26]. In contrast, higher angles of attack tend to increase the amount of thrust force in fluid environments [18, 27, 14]. A longer total screw length also leads to increased performance in water [18]. An in-depth study into screw-terrain interaction using terramechanics models determined that the ratio of blade height to screw radius is a vital factor for performance in soft granular media as it affects the level of sinkage and contact surface area, and that angle of attack should be chosen carefully to ensure that the screws can produce a positive tractive force when slipping [8]. In an analysis of screw performance in the sand across different parameter combinations for the angle of attack, blade height, and number of starts, optimal values were found to be 35°, 14 mm, and 1, respectively [28].
Outside of screw-based locomotion, reconfiguration for aerial locomotion has been explored in nature to understand how changing parameters of bird wings affect flight performance [29, 30]. Similar abilities have been replicated in robotics to achieve more efficient flight performance from wing designs [31]. Reconfigurable propeller designs have also been proposed to create more efficient propulsion in different scenarios [32, 33]. Similarly, NASU can dynamically change the angle of attack to adjust for different locomotion capabilities.
II System Design
The design of NASU, as noted previously, is structured after the origami kresling unit. The kresling unit defines a kinematic constraint between the length of the unit and the angle of the struts. By augmenting and controlling the total length of the unit, we can control the angle of attack. Its operation is shown in Figure 2. With our design, we verified the ability to produce appropriate thrust force and withstand reaction forces based on previous literature that captured data on screw forces. Finally, we integrated the NASU into a mobile testing platform which provided a means to gather performance across media.
Figure 2: Top and bottom rows show NASU at the minimum and maximum angles of attack. The angle of attack is set by linearly actuating the back plate whose position is denoted with parameter $d$. When the linear actuator moves, 2-DOF joints connecting the front and back plates to the struts are passively actuated to set the angle of attack of the blades. Finally, the angle of attack can be computed using $d_{0}$ and the strut length, denoted as $\ell$.
II-A Mechanical Design
@@
II-B NASU Locomotion Modeling and Metrics
@@
Figure 5: The mobile test bed was augmented to mount NASU for experimentation. The mobile test bed constrains the motion in a linear direction and measures the resultant propulsion forces and traveling velocity. The pulley system is used to apply a counterweight compensating for the test bed’s added mass which leaves the effective mass equal to that of NASU including its driving motor and linear actuator to adjust the angle of attack.
Figure 6: Using our previously developed mobile test bed that evaluates screw-based locomotion [24], we experiment over a large variety of media to validate NASU’s performance. The images from left to right are from the experiments conducted in the following media: small gravel, wet sand, big gravel, grass, mud, wood chips, and sand.
@@
II-C Mobile Test Bed: NASU Actuation
Modifications were made to the test bed to enable proper testing of NASU, and Figure 5 shows an overview of our changes to the test bed. First, a pulley system is implemented on the test bed to apply a counterweight that compensates for the additional mass not part of the NASU design (e.g. vertical bar, linear rail, 6-DOF FTS). The NASU unit by itself weighs 2.1kg, meanwhile, the entire mounted NASU system on the test bed weighs 6.2kg. Therefore, we apply 4.1kg of counterweight on the pulley, and this feature can be seen in Figure 5(b). Second, the motor housing for the linear actuator was secured to the vertical bar of the test bed as seen in Figure 5(c). A constraining plate is placed behind the back plate of NASU with its orientation constrained but linearly motion allowed. The ball screw drives the shortening or elongation of the mechanism through the movement of this plate. Finally, for the kresling unit design to properly morph to new angles of attack it was necessary to maintain a torsional degree of freedom which was accomplished through the addition of a lazy susan bearing sandwiched between the back plate of the NASU unit and the constraining plate driven by the linear actuator. The bearing is shown in Figure 5(d).
III Experiments and Results
@@ III-B Results
IV Discussion
As seen in the experimental results, NASU’s operating range for angle of attack sufficiently covers this trade-off. This trend for efficiency can be explained by several factors:
@@ IV Discussion
    As the angle of attack increases, the blades push more material to the sides and have more of an excavating effect, thus losing far more energy to the displacement of media.
@@
    For a given input torque, thrust force increases with decreasing angle of attack, compensating for the decrease in speed.
Figure 1 gives a simple diagram comparing the maximum and minimum angles of attack, as well as labeling the trade-off. This trade-off reinforces the vital importance of the angle of attack in the performance of screw-based locomotion and suggests that the angle of attack should be varied depending on the desired use case, highlighting the utility of NASU.
@@
Meanwhile, NASU was unable to produce significant propulsion in sand which is different from our previous results in screw-based locomotion [24]. This is likely due to two reasons. First, NASU lacks an internal shell that is found on conventional Archimedes screw propellers. The internal shell plays an important role of constraining the volume of material that is sheared by the screw blades so that the majority of the material is pushed along the axis of the screw. This effect is especially pronounced in softer media with lower shearing forces such as sand. Secondly, the NASU has shorter, straight blades as opposed to conventional helical blades, which would provide a more continuous shearing action. Figure 8 depicts the difference in the effectiveness of NASU in a compact media such as mud vs. a loose granular media such as sand.
We hypothesized that there would be a different optimal angle of attack, with respect to velocity or efficiency, for different media rather than the simple trade-off trend we found. Previous research in terramechanics screw-based propulsion suggests that the physical properties of different media should influence the performance [8]. Our results show that the velocity and efficiency trends are consistent across all media for the range of angles we tested. We believe that we found no environment-specific trends due to the specific range of angles we tested, differences in design of NASU compared to conventional screw propellers, or there simply is no environment-specific trends for screw-based locomotion.
@@
Figure 8: The image on the left shows NASU in mud, it leaves marks in the material as it provides a closer-to-solid surface to push off of when compressed, but is soft enough for the blades to initially pierce the media. The image on the right shows NASU in the sand, where the material is easy to initially pierce, but rather than compact, sand is instead dug out and pushed aside.
V Conclusion
@@
This work showcases the first mechanism to dynamically reconfigure the angle of attack of a screw propeller to optimize locomotion capability on the fly. We conduct experiments in a variety of real-world environments, demonstrating a trade-off exists between traveling velocity and locomotive efficiency in the range of NASU’s configuration space. Our intention with this mechanism is to enable future screw-based vehicles to control the angle of attack and make adjustments depending on their environment and use case. A higher angle of attack is more optimal when speed and mobility are a priority. While NASU was not tested in water or fluid environments, it is known that larger angles of attack, in the specific case of Archimedes-type screw propellers, produce more thrust and velocity in fluids [18, 27, 14], and we hypothesize that this trend will continue with NASU. Alternatively, a lower angle of attack is more optimal in solid environments if thrust force and efficiency are a priority, such as when towing a heavy load.
@@
In future work, the design may be improved by adding an internal shell as discussed in the previous section, and stacking multiple NASU units together to achieve a longer more continuous blade interaction. We also intend to experiment with different blade geometries through the replaceable blade connection. Another future work is to integrate this technology into ARCSnake [13, 14] or other screw-like robots to improve their versatility and adaptability.
VI Acknowledgements

### v3

# Calvin Joyce, Jason Lim, Roger Nguyen, Michael Owens, Sara Wickenhiser

@@ 1 Introduction
Figure 1: NASU is a screw-based propulsion unit for robot ground mobility that dynamically adjusts its angle of attack. The use of NASU is to allow control for optimizing efficiency or velocity over a wide array of terrains. It allows for trading off higher traveling velocity and lower efficiency (left) for higher efficiency and lower velocity (right) for resource-constrained robotic deployments.
@@ 1.2 Related Works
Some of the first studies in screw propulsion and locomotion showed theoretically and experimentally that velocity tends to increase with increasing angle of attack, while thrust force decreases in sand and soil-like environments [18, 19]. Later computational and experimental analyses confirmed a higher thrust force is produced with smaller angles of attack in the case of granular media composed of silica beads [25, 26]. In contrast, higher angles of attack tend to increase the amount of thrust force in fluid environments [18, 27, 14]. A longer total screw length also leads to increased performance in water [18]. An in-depth study into screw-terrain interaction using terramechanics models determined that the ratio of blade height to screw radius is a vital factor for performance in soft granular media as it affects the level of sinkage and contact surface area. That angle of attack should be chosen carefully to ensure that the screws can produce a positive tractive force when slipping [8]. In an analysis of screw performance in the sand across different parameter combinations for the angle of attack, blade height, and number of starts, optimal values were found to be 35°, 14 mm, and 1, respectively [28].
Outside of screw-based locomotion, reconfiguration for aerial locomotion has been explored in nature to understand how changing parameters of bird wings affect flight performance [29, 30]. Similar abilities have been replicated in robotics to achieve more efficient flight performance from wing designs [31]. Reconfigurable propeller designs have also been proposed to create more efficient propulsion in different scenarios [32, 33]. Adjustable diameter wheels have been proposed to be able to climb over different heights of obstacles while still being able to get into small places for search and rescue applications [34]. Similarly, NASU can dynamically change the angle of attack to adjust for different locomotion capabilities.
2 System Design
The design of NASU, as noted previously, is structured after the origami kresling unit. The kresling unit defines a kinematic constraint between the length of the unit and the angle of the struts. By augmenting and controlling the total length of the unit, we can control the angle of attack. Its operation is shown in Figure 2. With our design, we verified the ability to produce appropriate thrust force and withstand reaction forces based on previous literature that captured data on screw forces. Finally, we integrated the NASU into a mobile testing platform which provided a means to measure performance across media.
Figure 2: From left to right, the figures show NASU at the minimum and maximum angles of attack. The angle of attack is set by linearly actuating the back plate whose position is denoted with parameter $d$. When the linear actuator moves, 2-DOF joints connecting the front and back plates to the struts are passively actuated to set the angle of attack of the blades. Finally, the angle of attack can be computed using $d_{0}$ and the strut length, denoted as $\ell$.
2.1 Mechanical Design
@@
2.2 NASU Locomotion Modeling and Metrics
@@
Figure 5: The mobile test bed was augmented to mount NASU for experimentation. The mobile test bed constrains the motion linearly and measures the resultant propulsion forces and traveling velocity. The pulley system is used to apply a counterweight compensating for the test bed’s added mass which leaves the effective mass equal to that of NASU including its driving motor and linear actuator to adjust the angle of attack.
2.3 Mobile Test Bed: NASU Actuation
Modifications were made to the test bed to enable proper testing of NASU, and Figure 5 shows an overview of our changes to the test bed. First, a pulley system is implemented on the test bed to apply a counterweight that compensates for the additional mass not part of the NASU design (e.g. vertical bar, linear rail, 6-DOF FTS). The NASU unit by itself weighs 2.1kg; meanwhile, the entire mounted NASU system on the test bed weighs 6.2kg. Therefore, we apply 4.1kg of counterweight on the pulley, and this feature can be seen in Figure 5(b). Second, the motor housing for the linear actuator is secured to the vertical bar of the test bed as seen in Figure 5(c). A constraining plate is placed behind the back plate of NASU with its orientation constrained but linearly motion allowed. The ball screw drives the shortening or elongation of the mechanism through the movement of this plate. Finally, for the kresling unit design to properly morph to new angles of attack it was necessary to maintain a torsional degree of freedom which was accomplished through the addition of a lazy susan bearing sandwiched between the back plate of the NASU unit and the constraining plate driven by the linear actuator. The bearing is shown in Figure 5(d).
Figure 6: Using our previously developed mobile test bed that evaluates screw-based locomotion [24], we experiment over various media to validate NASU’s performance. The images from left to right are from the experiments conducted in the following media: small gravel, wet sand, big gravel, grass, mud, wood chips, and sand.
@@ 2.3 Mobile Test Bed: NASU Actuation
3 Experiments and Results
@@ 3.2 Results
4 Discussion
As seen in the experimental results, NASU’s operating range for angle of attack, 10-35, provides a trade-off between locomotion velocity and efficiency. While we anticipated a positive correlation between the angle of attack and forward velocity based on previous research in screw-based locomotion [24, 18, 19], the negative correlation with efficiency was a finding from our experimental results. We believe the efficiency trend can be explained by several factors:
@@ 4 Discussion
    As the angle of attack increases, the relative motion of the material increases meaning the material is moved backward faster, driving the increase in velocity. However, the blades push more material to the sides, excavating rather than contributing to thrust, thus losing far more energy to the displacement of media.
@@
    For a given input torque, the available thrust force increases with decreasing angle of attack compensating for the decrease in speed. This means the towing capacity of the robot increases at the cost of moving quickly.
Figure 1 gives a simple diagram comparing the maximum and minimum angles of attack, as well as labeling the trade-off. This trade-off reinforces the importance of the angle of attack in the performance of screw-based locomotion and suggests that the angle of attack should be varied depending on the desired use case, highlighting the future utility of NASU.
@@
Figure 8: The image on the left shows NASU in mud, it leaves marks in the material as it provides a closer-to-solid surface to push off of when compressed, but is soft enough for the blades to pierce the media initially. The image on the right shows NASU in the sand, where the material is easy to pierce initially, but rather than compact, sand is instead dug out and pushed aside.
Meanwhile, NASU was unable to produce significant propulsion in sand which is different from our previous results in screw-based locomotion [24]. This is likely due to two reasons. First, NASU lacks an internal shell that is found on conventional Archimedes screw propellers. The internal shell plays an important role in constraining the volume of material that is sheared by the screw blades so that the majority of the material is pushed along the axis of the screw. This effect is especially pronounced in softer media with lower shearing forces such as sand. Secondly, the NASU has shorter, straight blades as opposed to conventional helical blades, which would provide a more continuous shearing action. Figure 8 depicts the difference in the effectiveness of NASU in a compact media such as mud vs. a loose granular media such as sand.
We hypothesized that there is a different optimal angle of attack, concerning velocity or efficiency, for different media rather than the simple trade-off trend we found. Previous research in terramechanics screw-based propulsion suggests that the physical properties of different media should influence the performance [8]. Our results show that the velocity and efficiency trends are consistent across all media for the range of angles we tested. We believe that we found no environment-specific trends due to the specific range of angles we tested, differences in the design of NASU compared to conventional screw propellers, or there simply are no environment-specific trends for screw-based locomotion.
5 Conclusion
@@
This work showcases the first mechanism to dynamically reconfigure the angle of attack of a screw propeller to optimize locomotion capability on the fly. We conduct experiments in a variety of real-world environments, demonstrating a trade-off exists between traveling velocity and locomotive efficiency in the range of NASU’s configuration space. Our intention with this mechanism is to enable future screw-based vehicles to control the angle of attack and make adjustments depending on their environment and use case. For example, a higher angle of attack is more optimal when speed and mobility are a priority as anticipated from previous literature [24, 18, 19]. Alternatively, a key finding of our experiments is that a lower angle of attack is preferred if thrust force and efficiency are a priority, such as when towing a heavy load. While NASU was not tested in water or fluid environments, it is known that larger angles of attack, in the specific case of Archimedes-type screw propellers, produce more thrust and velocity in fluids [18, 27, 14], and we hypothesize that this trend will continue with NASU.
@@
In future work, the design may be improved by adding an internal shell as discussed in the previous section, and stacking multiple NASU units together to achieve a longer more continuous blade interaction. We also intend to experiment with different blade geometries through the replaceable blade connection. Another future work is to integrate this technology into ARCSnake [13, 14] or other screw-based robots to improve their versatility and adaptability.
6 Acknowledgements